OpenStudy (anonymous):

Simplify ((1/y-b)-(1/y))/b

OpenStudy (anonymous):

\[\frac{ (\frac{ 1 }{ y-b }) - (\frac{ 1 }{ y }) }{ b }\] it looks like this, isn't it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\frac{ (\frac{ 1 }{y-b }) - (\frac{ 1 }{ y }) }{ b } = \frac{ \frac{ y -(y-b) }{ y(y-b) } }{ b } =\frac{ \frac{ y-y+b }{ y(y-b) } }{ b } = \frac{ \frac{ b }{y(y-b) } }{ b } = \frac{ b }{ y(y-b) } \times \frac{ 1 }{ b } = \frac{ 1 }{ y(y-b) }\]

OpenStudy (anonymous):

get it ??

OpenStudy (anonymous):

So \[\frac{ 1 }{ y(y-b) }\]

OpenStudy (anonymous):

so for the reciprocal of b you don't multiply by b/1 right it's 1/b?

OpenStudy (anonymous):

@gerryliyana ?

OpenStudy (anonymous):

\[\frac{ \frac{ b }{ y(y-b) } }{ b } = \frac{ \frac{ b }{ y(y-b) } }{ \frac{ b }{ 1 } } = \frac{ b }{ y(y-b) } \times \frac{ 1 }{ b }\]

OpenStudy (anonymous):

oh i see

OpenStudy (anonymous):

get it?

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

yeah thank you!

OpenStudy (anonymous):

you are very welcome :)